A wave is described by y = 0.020 8 sin(kx - wt), where k = 2.22 rad/m, w = 3.66 rad/s, x and y are in meters, and t is in seconds

6. A wave is described by y = 0.020 8 sin(kx - wt), where k = 2.22 rad/m, w = 3.66 rad/s, x and y are in meters, and t is in seconds. 

(a) Determine the amplitude of the wave. 

(b) Determine the wavelength of the wave. 

(c) Determine the frequency of the wave. 

(d) Determine the speed of the wave. 


7. When a particular wire is vibrating with a frequency of 3.00 Hz, a transverse wave of wavelength 64.0 cm is produced. Determine the speed of waves along the wire. 


8. Transverse waves travel with a speed of 18.0 m/s on a string under a tension of 6.20 N. What tension is required for a wave speed of 28.0 m/s on the same string? 
Answer #1

1a) amplitude = 0.0208 m

b) wavelength = 2pi/k = 2pi/2.22 = 2.83 m

c] frequency = w/2pi = 3.66/2pi = 0.5825 Hz

d] speed = wavelength*frequency = 2.83*0.5825 = 1.65 m/s